# CompBench: A Comparative Reasoning Benchmark for Multimodal LLMs

### v1

@@ A.1 Limitations
While we conducted a human evaluation study to establish the upper bound performance on CompBench, the study is currently limited to 140 samples assessed by five evaluators (cf. §5.3 in the main text). We plan to expand the study to a larger scale in future work.
@@ A.2 Social impacts
CompBench evaluates the comparative reasoning abilities of MLLMs in images. A potential negative impact of our work is that malicious users might exploit our concept (i.e., comparison) to compare ethical or offensive content. Therefore, it is essential to incorporate effective safeguards in MLLMs to filter out any inappropriate materials.
@@ A.3 Ethical considerations
All fourteen datasets (cf. Table 1 in the main text) that we used to curate CompBench adhere to strict guidelines to exclude any harmful, unethical, or offensive content. Additionally, we instruct human annotators to avoid generating any personally identifiable information or offensive content during our annotation process. Finally, we do not conduct any study to compare harmful, ethical, or offensive content between the two images.
@@ A.4 License of assets
All fourteen datasets are publicly available, and Table 5 details the licensing information for the assets in each dataset. We release our CompBench under a Creative Commons Attribution 4.0 License (CC BY 4.0) to enhance global accessibility and foster innovation and collaboration in research.
@@
Appendix B CompBench Curation Details
@@ B.1 Annotation Details
MagicBrush [61] is a large-scale, manually annotated dataset for instruction-guided real image editing. For each image, MagicBrush utilizes DALL-E 2 [48] to generate an edited version of the image based on language instructions, such as “let the flowers in the vase be blue.” Our goal is to identify pairs of similar images. We thus use CLIP [46] to evaluate the visual similarity between the original and edited images. Only pairs exceeding a predetermined similarity threshold are selected as candidate samples for our CompBench. For each selected pair, we then construct a multiple-choice question to ask the difference between two images in the pairs. Concretely, we first use GPT-4V [1] to extract all relevant objects and their attributes from the edited image with the following prompt:
@@
SoccerNet [19], CompCars [59], VQAv2 [21], Q-bench2 [62] are automatically processed to generate samples for CompBench using their metadata and CLIP visual similarity. For more details, please refer to §4.2 of the main text.

### v2

@@ A.1 Limitations
While we conducted a human evaluation study to establish the upper bound performance on MLLM-CompBench, the study is currently limited to 140 samples assessed by five evaluators (cf. §5.3 in the main text). We plan to expand the study to a larger scale in future work.
@@ A.2 Social impacts
MLLM-CompBench evaluates the comparative reasoning abilities of MLLMs in images. A potential negative impact of our work is that malicious users might exploit our concept (i.e., comparison) to compare ethical or offensive content. Therefore, it is essential to incorporate effective safeguards in MLLMs to filter out any inappropriate materials.
@@ A.3 Ethical considerations
All fourteen datasets (cf. Table 1 in the main text) that we used to curate MLLM-CompBench adhere to strict guidelines to exclude any harmful, unethical, or offensive content. Additionally, we instruct human annotators to avoid generating any personally identifiable information or offensive content during our annotation process. Finally, we do not conduct any study to compare harmful, ethical, or offensive content between the two images.
@@ A.4 License of assets
All fourteen datasets are publicly available, and Table 6 details the licensing information for the assets in each dataset. We release our MLLM-CompBench under a Creative Commons Attribution 4.0 License (CC BY 4.0) to enhance global accessibility and foster innovation and collaboration in research.
@@
Appendix B MLLM-CompBench Curation Details
@@ B.1 Annotation Details
MagicBrush [66] is a large-scale, manually annotated dataset for instruction-guided real image editing. For each image, MagicBrush utilizes DALL-E 2 [51] to generate an edited version of the image based on language instructions, such as “let the flowers in the vase be blue.” Our goal is to identify pairs of similar images. We thus use CLIP [49] to evaluate the visual similarity between the original and edited images. Only pairs exceeding a predetermined similarity threshold are selected as candidate samples for our MLLM-CompBench. For each selected pair, we then construct a multiple-choice question to ask the difference between two images in the pairs. Concretely, we first use GPT-4V [1] to extract all relevant objects and their attributes from the edited image with the following prompt:
@@
SoccerNet [19], CompCars [64], VQAv2 [21], Q-bench2 [67] are automatically processed to generate samples for MLLM-CompBench using their metadata and CLIP visual similarity. For more details, please refer to §4.2 of the main text.
@@ Appendix D More qualitative examples
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See discussions in Suppl.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See discussions in Suppl.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See details in Suppl.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See details in Suppl.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Most of our experiments are evaluating existing MLLM models.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See details in Suppl.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] See details in Suppl.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We curated new annotations.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See discussions in Suppl.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See discussions in Suppl.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]